OpenStudy (anonymous):

8x = 4 - x = 6x = 5 = x solve and state if the problem is all reall numbers or a contradiction (no solution) show work

OpenStudy (anonymous):

x=5 then 8x is not equal to 6x

OpenStudy (anonymous):

I mean 8x + 4 - x = 6x + 5 + x

OpenStudy (agreene):

8x + 4 - x = 6x + 5 + x 7x + 4 = 7x + 5 <== cannot be true for real numbers